OpenStudy (anonymous):

Please help me find the y-intercept 5x-2y=4y+3x+7

OpenStudy (anonymous):

The y-intercept occurs when x = 0 5x-2y=4y+3x+7 5(0)-2y=4y+3(0)+7 0-2y=4y+0+7 0-2y-4y=7 0-6y=7 y=7/6

OpenStudy (anonymous):

would that be 0,-7/6 or -7/6,0

OpenStudy (anonymous):

What? :/

OpenStudy (anonymous):

-7/6 will be y-intercept and x-intercept will be 0